OpenStudy (anonymous):

during a rain storm the level of a river rose 1 1/4 inches in 4 1/2 hours. how many inches did it rise in one hour

OpenStudy (anonymous):

\[v=x/t=11/82\] \[x = vt = 1 * 11/8= 11/8 inch\]